# GRIP: Algorithm-Agnostic Machine Unlearning for Mixture-of-Experts via Geometric Router Constraints

### v1

@@ C.1 Problem Formulation
C.2 Algorithm
Algorithm 1 details the procedure. We pre-compute the row norms and sampling probabilities. At each step, we sample a constraint; if the constraint is violated, we project the current gradient estimate orthogonally onto the defining hyperplane.
@@ C.4 Computational Complexity
C.5 Complexity and Memory Analysis
@@
Appendix D Ablation Studies
D.1 Threshold Analysis
In comparing the threshold values $\epsilon$, we tested values across four orders of magnitude $(\epsilon\in\{10^{-4},10^{-3},10^{-2},10^{-1}\})$. We use $\epsilon=10^{-2}$ for all experiments. Values orders of magnitude less than $10^{-2}$ proved overly restricted and resulted in empty null spaces in later layers. Detailed analysis is contained in Appendix D. Below contains an analysis of valid thresholds that allow all null spaces to exist.
Table 6: Sensitivity Analysis of Null-Space Threshold $\epsilon$
Threshold ($\epsilon$) RS FA RA
$10^{-4}$ 0.97 0.31 0.59
$10^{-3}$ 0.96 0.28 0.62
$10^{-2}$ (Default) 0.96 0.24 0.65
$10^{-1}$ 0.89 0.25 0.59

### v2

@@ D.1 Problem Formulation
D.2 Algorithm
Algorithm 1 details the procedure. We precompute the row norms and sampling probabilities. At each step, we sample a constraint; if the constraint is violated, we project the current gradient estimate orthogonally onto the defining hyperplane.
@@ D.4 Computational Complexity
D.5 Complexity and Memory Analysis
@@
Appendix E Security Evaluation Protocols
To rigorously quantify the robustness of unlearning, we define two formal evaluation algorithms: Adversarial Expert Forcing, which tests for latent knowledge retention in expert parameters, and Side-Channel Vulnerability, which measures information leakage through routing patterns.
E.1 Adversarial Expert Forcing
This protocol simulates a white-box adversary who bypasses the router’s suppression mechanism to directly access erased knowledge. We define the recovery rate as the accuracy of the unlearned model when forced to use the routing path of the original, pre-unlearning model.
1:Input: Pre-unlearning Model $\Theta_{pre}$, Unlearned Model $\Theta_{post}$, Hazardous Query Set $\mathcal{D}_{haz}$
2:Output: Adversarial Recovery Rate (ARR)
[…]
4:for each input pair $(x,y)$ in $\mathcal{D}_{haz}$ do
5:  {Step 1: Extract original routing path}
6:  $\mathcal{S}_{orig}\leftarrow\text{GetActiveExperts}(\Theta_{pre},x)$
7:  {Step 2: Forward pass with forced routing}
8:  {Override $\Theta_{post}$ router to strictly use $\mathcal{S}_{orig}$}
9:  $\hat{y}\leftarrow\text{Forward}(\Theta_{post},x,\text{force\_experts}=\mathcal{S}_{orig})$
10:  {Step 3: Measure recovery}
11:  if $\text{argmax}(\hat{y})=y$ then
12:   $N_{correct}\leftarrow N_{correct}+1$
13:  end if
14:end for
15:return $N_{correct}/|\mathcal{D}_{haz}|$
Algorithm 2 Adversarial Expert Forcing Evaluation
E.2 Side-Channel Vulnerability Assessment
This protocol quantifies the distinguishability of “forgotten” queries from retained queries based solely on expert activation patterns. This simulates a side-channel attack where an adversary infers query membership via timing or power analysis of specific GPU nodes in a distributed inference setting.
1:Input: Unlearned Model $\Theta_{post}$, Forget Set $\mathcal{D}_{f}$, Retain Set $\mathcal{D}_{r}$
2:Output: Vulnerability Ratio (VR)
3: {Step 1: Construct Fingerprint Dataset}
[…]
5:for each $x\in\mathcal{D}_{f}\cup\mathcal{D}_{r}$ do
6:  $\mathcal{S}_{x}\leftarrow\text{GetActiveExperts}(\Theta_{post},x)$
7:  $v_{fp}\leftarrow\mathbf{0}\in\{0,1\}^{E}$ {Initialize binary fingerprint vector}
8:  for each expert index $e\in\mathcal{S}_{x}$ do
9:   $v_{fp}[e]\leftarrow 1$
10:  end for
11:  $y_{label}\leftarrow 1\text{ if }x\in\mathcal{D}_{f}\text{ else }0$
12:  $\mathcal{X}_{probe}.\text{append}((v_{fp},y_{label}))$
13:end for
14: {Step 2: Train Attack Classifier (Logistic Regression)}
15: Split $\mathcal{X}_{probe}$ into $\mathcal{X}_{train},\mathcal{X}_{test}$
16: Fit classifier $\mathcal{C}$ on $\mathcal{X}_{train}$ to predict $y_{label}$ from $v_{fp}$
17: {Step 3: Evaluate Leakage}
[…]
20:return $\text{Accuracy}/\text{RandomBaseline}$ {VR >1.0>1.0 implies leakage}
Algorithm 3 Side-Channel Leakage Quantification
Appendix F Direct Comparison with SEUF
To evaluate the efficacy of our framework against the current state-of-the-art in MoE unlearning, we perform a direct head-to-head comparison with SEUF (Sparse Expert Unlearning Framework) [43].
Experimental Setup.
To ensure a strictly fair comparison, we replicate the specific experimental conditions used in the original SEUF study. We utilize the Qwen1.5-MoE-A2.7B-Chat architecture instead of the 30B model used in our main experiments. We evaluate performance on the WMDP-Cyber dataset, consistent with the hazardous knowledge removal tasks targeted by both frameworks. We compare both methods when applied to the Representation Misdirection (RMU) unlearning algorithm.
Table 7: Direct comparison with SEUF on Qwen1.5-MoE-A2.7B-Chat. $\downarrow$$\uparrow$ indicates lower/higher is better. Our results are bolded.
[…]
RMU Baseline 0.28 0.38 0.35
SEUF + RMU 0.25 0.53 0.87
Relaxed Constraint (GRIP) + RMU 0.25 0.61 0.94
PTC (GRIP) + RMU 0.27 0.63 0.99
F.1 Results and Analysis
The quantitative results are presented in Table 7. Our analysis highlights three key findings:
Unlearning Efficacy (FA):
Both frameworks successfully reduce the model’s accuracy on the forget set to near-random levels. GRIP (Relaxed Constraint) achieves an FA of 0.25, matching SEUF exactly, while PTC achieves 0.27. This confirms that restricting the router’s optimization landscape via null-space projections does not hinder the model’s ability to erase hazardous knowledge.
Superior Utility Preservation (RA):
GRIP demonstrates a significant advantage in retaining general model capabilities. While SEUF recovers a Retain Accuracy (RA) of 0.53, GRIP improves this to 0.61 (+15%) with relaxed constraints and 0.63 (+19%) with PTC. SEUF relies on soft regularization penalties ($L_{2}$ norms on router weights) to maintain stability. Our results suggest these soft penalties are insufficient to prevent the "optimization shortcut" described in Section 1, leading to degraded utility. In contrast, GRIP’s hard geometric constraints strictly enforce the preservation of retain-set routing, protecting general knowledge more effectively.
Routing Stability (RS):
We observe that SEUF improves routing stability over the unconstrained baseline (0.87 vs 0.35) but still permits noticeable drift. GRIP eliminates this drift almost entirely, achieving an RS of 0.94 with training-time constraints and 0.99 with Post-Training Correction. This validates our hypothesis that decoupling routing stability from parameter plasticity is essential for robust MoE unlearning.